# Simulating Dual-Pixel Images From Ray Tracing For Depth Estimation

### v1

@@ Abstract
Many studies utilize dual-pixel (DP) sensor phase characteristics for various applications, such as depth estimation and deblurring. However, since the DP image features are entirely determined by the camera hardware, DP-depth paired datasets are very scarce, especially when performing depth estimation on customized cameras. To overcome this, studies simulate DP images using ideal optical system models. However, these simulations often violate real optical propagation laws, leading to poor generalization to real DP data. To address this, we investigate the domain gap between simulated and real DP data, and propose solutions using the Simulating DP images from ray tracing (Sdirt) scheme. The Sdirt generates realistic DP images via ray tracing and integrates them into the depth estimation training pipeline. Experimental results show that models trained with Sdirt-simulated images generalize better to real DP data. The code and collected datasets will be available on https://github.com/LinYark/Sdirt.
@@ 1 Introduction
The DP sensor [15, 29] is designed to split each pixel into a left and a right sub-pixel, utilizing microlenses and sub-pixels to achieve phase splitting. This allows capturing a pair of images in one shot, known as DP images, as shown in Fig. 1(a). DP images can not only be used to autofocus [31, 11, 7] but also enhance the performance in applications such as deblurring [41, 17, 4] and depth estimation [13, 9, 23]. In the field of Depth-from-Dual-Pixel (DfDP) depth estimation, DP-depth paired datasets are scarce due to camera hardware limitations. To address the scarcity of DP-depth paired datasets, much research [37, 26, 18, 23, 2, 16] has focused on simulating DP images using RGBD datasets, aiming to provide a more flexible and accessible data source.
@@
Figure 1: (a) Imaging process of a DP camera. We illustrate the slight shifts in the left and right DP images caused by phase splitting with white dashed lines. (b) Example comparison between real and COC simulated DP PSF. There exists a significant difference between the two.
@@
The key to simulating the DP image through RGBD datasets lies in the DP Point-Spread-Function (PSF). In recent years, many model-based and calibration-based DP PSF simulators have emerged. Calibration-based simulators [37, 16, 18] spent considerable time calibrating real cameras to match the DP PSF but faced issues such as interpolation errors from discrete calibration and neglect of lens and DP sensor parameters, limiting model transferability. Model-based simulators [26, 23, 2] used ideal optical models or aberration analysis models to directly calculate the DP PSF, thus bypassing the calibration step. However, these models perform poorly on real DP images due to ignoring lens aberration and DP phase splitting characteristics. As shown in Fig. 1(b), the ideal thin lens model’s simulated Circle-of-Confusion (COC) DP PSF has a significant domain gap with the real one, affecting depth estimation.
@@
To address the domain gap between real DP images and DP images simulated by model-based simulators, we propose the Sdirt scheme. Sdirt consists of a ray-traced DP PSF simulator and a pixel-wise DP image rendering module, which can accurately generate DP images with aberration and phase characteristics. Specifically, for a fixed focus lens and DP sensor system with known parameters, we use ray tracing [35, 40, 39] to calculate the spatially varying DP PSF and train a multilayer perceptron (MLP) to predict it. Subsequently, we convolve the per-pixel DP PSF with the All-in-Foucs (AiF) image to simulate the DP images captured by a real camera. After training a DfDP model with DP images simulated by Sdirt, the model can accurately estimate depth using optical aberration and phase information, which enhances its ability to generalize to real DP images.
@@
We conduct extensive experiments to validate the effectiveness of Sdirt. The experimental results indicate that, compared to other model-based simulators, the DP PSFs and DP images generated by Sdirt are closer to real images, and the DfDP model trained on Sdirt achieves better generalization. In summary, our contributions are four-fold:
@@
  • We propose a ray-traced DP PSF simulator that calculates the spatially varying DP PSF, addressing the domain gap between simulated and real DP PSF caused by lens aberrations and sensor phase splitting.
@@
  • Experimental results of depth estimation demonstrate that the DfDP model trained on Sdirt generalizes better to real DP images.
  • We collected a DP-depth paired real test set, with an open lens structure and fixed focus, for use in subsequent research.
@@ 2.1 DP Datasets
Most professional and mobile cameras have DP sensors, but only Canon cameras [1, 25, 26] and Google Pixel phones [43, 9, 34] provide DP data. In the field of depth estimation, supervised tasks rely on DP-depth paired datasets [9, 43, 26, 1]. However, when the lens or sensor is changed, or the focal length and aperture settings are adjusted, the point spread function (PSF) is altered, requiring significant time and effort to re-collect DP-depth paired data. Consequently, many researchers are focusing on simulating DP-depth paired datasets using RGBD datasets.
2.2 DP PSF simulator
Simulating DP images from RGBD datasets relies on DP PSF, with recent progress in model-based [37, 18, 16] and calibration-based [26, 23, 2] DP PSF simulators. In calibration-based simulators, Xin et al. [37] acquired full-space DP PSF by sampling discrete points in space and interpolating, while Li et al. [18] acquired it through size scaling using the COC model. Furthermore, Li et al. [16] generated DP PSF through a U-Net network, which requires a large amount of DP-depth paired data. Calibration-based simulators are time-consuming and prone to interpolation and scaling errors, while model-based simulators circumvent these issues. Pan et al. [23] simplified the optical system as an ideal thin lens and used a symmetrically divided rectangular aperture to calculate DP PSF. Punnappurath et al. [26] fitted DP PSF as the left-right flip symmetric kernel depletion shape. Abuolaim et al. [2] used aberration analysis models to calculate DP PSF, but the resulting DP PSF remained symmetric. These studies oversimplify optical propagation, neglecting aberrations and phase splitting, causing a large gap between simulated and real DP data.
@@ 2.3 Applications driven by DP data
In recent years, DP data-driven applications such as depth estimation [13, 23, 9, 43, 14, 24, 10], deblurring [41, 17, 4, 38, 3, 42, 27], and refocus [11, 7, 32, 5] have undergone rapid development. Additionally, Kang et al. [13] utilized DP data for facial scanning, similar to depth estimation tasks. Wadhwa et al. [34] employed DP data for synthetic shallow depth-of-field imaging, a popular application in mobile photography recently. Shi et al. [28] integrated DP hardware with structured diffractive optical elements for complex image encoding and high-precision multimodal reconstruction. In addition, DP data can also be used for disparity estimation [36, 22], rain [18], and reflections [25] removal. It is foreseeable that DP data-driven applications will continue to emerge in various new fields.
Figure 2: Simulating dual pixel images from ray tracing pipeline. (a) Ray-traced DP PSF simulator. Calculates spatially varying DP PSF for lens and DP sensor through ray tracing. (b) DP PSF prediction network. Trains an MLP network to predict DP PSF, using ray-traced DP PSF as ground truth. (c) Pixel-wise DP image rendering module. The network predicts the DP PSFs for all points in the depth map (red pass). Then, the DP PSF is convolved with the AiF RGB image to render the simulated DP image (blue pass).
@@ 3 Method of Sdirt
Training the DfDP network requires DP images as input. As shown in Fig. 2(c), the module achieves pixel-wise rendering of DP images by convolving the DP PSF corresponding to each point in the depth map with AiF RGB image. The DP PSF for each point is obtained through inference using a pre-trained MLP network. As illustrated in Fig. 2(b), the offline training process of the MLP network uses ray-traced DP PSF (Fig. 2(a)) as Ground Truth (GT).
Figure 3: (a) DP composite pixel structure. The left side is a right view, and the right side is a front perspective view. (b) On hitting the microlens, the kth ray lands on left/right sub-pixel in red/blue interval, else missed. (c) Without hitting the microlens, the kth ray lands on left/right sub-pixel in red/blue interval, else missed.
@@ 3.1 Ray-traced DP PSF simulator
As shown in Fig. 2(a), lens ray tracing can accurately obtain the landing point on the sensor surface [35, 40, 39, 8]. We construct a ray set $A$ containing $n$ rays, with all rays originating from a spatial object point $p$. Then, the image of the aperture stop in the object space of the lens is regarded as the entrance pupil, and $n$ points are densely sampled on this entrance pupil plane. The initial direction vector of each ray in $A$ is set to point from $p$ to the corresponding sampling point of the entrance pupil sequentially. As $A$ traverses each surface of the lens, its position and direction strictly follow the unique optical characteristics of the mirror surface, undergoing precise refraction according to Snell’s law. After a series of refractions through the lens, $A$ finally lands on the sensor plane. We denote $A$’s landing point on the sensor as $O$ and its ray direction as $D$.
DP sensor ray tracing is limited by camera manufacturers’ non-disclosure of microlens and sub-pixel structure. As a result, we simplify the DP composite pixel structure (Fig. 3(a)) based on past research [15, 29]. We model the microlens as a thin lens with radius $r$ and focal length $f$, setting $h$ for distance between the sub-pixel and the microlens, $w$ for sub-pixel width, and $ps$ for DP composite pixel size.
The DP PSF can be obtained by separately calculating the cumulative integration of rays for each sub-pixel. We analyze which sub-pixel the kth ray $A_{k}$ in the ray set $A$ ultimately enters. Through lens ray tracing, we know the landing point $O_{k}(x_{k},y_{k},z_{k})$ and direction $D_{k}(x_{k}^{\prime},y_{k}^{\prime},z_{k}^{\prime})$ of the kth ray $A_{k}$ on the sensor surface. Based on the landing point $O_{k}$, we can easily calculate that $A_{k}$ lies within the DP composite pixel $(i,j)$, having surface center coordinates $(x_{i},y_{i})$. For $O_{k}$, the situations differ depending on whether it is within or outside the microlens:
When $O_{k}$ is within the microlens, as shown in Fig. 3(b), $A_{k}$ is refracted by the microlens and then enters a sub-pixel. According to geometric optics theory, $A_{k}$ and the principal ray incident on the thin lens at the same angle converge on the focal plane of the thin lens. We only need to focus on a few boundary lines $x_{L1},x_{M1},x_{R1}$ to determine which sub-pixel $A_{k}$ currently enters. If $x_{k}$ is in the interval $[x_{L1},x_{M1}]$, it enters the left sub-pixel; if in the interval $[x_{M1},x_{R1}]$, it enters the right sub-pixel; otherwise, it is considered a missing ray. The calculation methods for the boundary lines are as follows:
@@
Where $\tan\theta$ is the tangent angle at which $A_{k}$ landing on the sensor surface, equivalent to ${x_{k}^{\prime}}/{z_{k}^{\prime}}$.
When $O_{k}$ is outside the microlens, as shown in Fig. 3(c), $A_{k}$ enters the sub-pixel directly without refraction. Similarly, we only need to focus on a few boundary lines $x_{L2},x_{M2},x_{R2}$ to determine which sub-pixel $A_{k}$ currently enters:
@@
The left PSF $PSF_{L}$ can be calculated by integrating the set of rays $A$ emitted from point $p$ over all left sub-pixels:
@@
where $\delta_{L,k}(i,j)$ represents the energy distribution of ray $A_{k}$ in the left sub-pixel at DP composite pixel $(i,j)$. The ray energy distribution is assumed to be a unit impulse, which equals zero unless it eventually enters the left sub-pixel of DP composite pixel $(i,j)$. By following the same calculation steps as for $PSF_{L}$, $PSF_{R}$ can also be computed.
Figure 4: (a) The valid imaging region is a frustum, and we normalize the $(x,y)$ coordinates to $[-1,1]$. (b) The DfDP model takes DP image as input to predict the detph map. (c) During cost volume generation in [6], we stack original disparity (green arrows) and add reverse disparity (blue arrows), with $dmax$ as max displacement.
@@ 3.2 Pixel-wise DP image rendering module
As shown in Fig. 2(b), we reduce computational costs of ray-traced DP PSF by training an MLP network to predict it, inspired by [33, 39]. The network input is the normalized coordinate $p$ within the camera’s valid imaging region (Fig. 4(a)), which is a frustum defined by field of view, sensor size, and the set minimum and maximum depths. We normalize the $(x,y)$ coordinates to $[-1,1]$. After fixing the focus distance, the kernel size $ks$ of the DP PSF is set to a size such that the DP PSF for any point within the valid imaging region can be fully displayed. Finally, we configure the network structure with 5 hidden layers, each containing 512 neurons, and an output layer with $2*ks^{2}$ neurons. We use the L2 function as the loss:
@@
During training, we use ray-traced DP PSFs as GT and apply max normalization to avoid challenging samples. During inference, we apply sum normalization to the predicted DP PSFs to mimic the vignetting correction in the camera image signal processor.
Due to the lack of shift-invariance of DP PSF, we perform local convolution of each pixel’s DP PSF with the AiF image to simulate the DP image captured by a camera. The RGBD dataset [30, 21] provides the paired RGB image $I_{RGB}$ and depth map $I_{D}$. As shown in Fig. 2(c), we treat each pixel in $I_{D}$ as a spatial object point and use the trained MLP network to predict its DP PSF. Using $I_{RGB}$ as the AiF image, we perform pixel-wise local convolution [39] to render DP image containing aberration and phase information.
Figure 5: DP PSF simulator representation result. Evaluate the real and simulated (ours, COC, L2R [2], Modeling [26] and DDDNet [23] ) F/1.8 DP PSFs at two depths (0.5m, 1.5m) and three different positions. As spatial object point $p$ moves further from the optical axis, the real $PSF_{L}$ and $PSF_{R}$ become more asymmetric, and aberrations increase. Their simulators neglect aberrations and DP phase splitting, causing a large gap between simulated and real DP PSF. Only our ray-traced simulator predicts realistic results at all depths and positions.

### v2

@@ Abstract
Many studies utilize dual-pixel (DP) sensor phase information for various applications, such as depth estimation and deblurring. However, since DP image features are entirely determined by the camera hardware, DP-depth paired datasets are very scarce, especially when performing depth estimation on customized cameras. To overcome this, studies simulate DP images using ideal optical models. However, these simulations often violate real optical propagation laws, leading to poor generalization to real DP data. To address this, we investigate the domain gap between simulated and real DP data, and propose solutions using the Simulating DP Images from Ray Tracing (Sdirt) scheme. Sdirt generates realistic DP images via ray tracing and integrates them into the depth estimation training pipeline. Experimental results show that models trained with Sdirt-simulated images generalize better to real DP data. The code and collected datasets will be available at https://github.com/LinYark/Sdirt.
Abstract
Many studies utilize dual-pixel (DP) sensor phase information for various applications, such as depth estimation and deblurring. However, since DP image features are entirely determined by the camera hardware, DP-depth paired datasets are very scarce, especially when performing depth estimation on customized cameras. To overcome this, studies simulate DP images using ideal optical models. However, these simulations often violate real optical propagation laws, leading to poor generalization to real DP data. To address this, we investigate the domain gap between simulated and real DP data, and propose solutions using the Simulating DP Images from Ray Tracing (Sdirt) scheme. Sdirt generates realistic DP images via ray tracing and integrates them into the depth estimation training pipeline. Experimental results show that models trained with Sdirt-simulated images generalize better to real DP data. The code and collected datasets will be available at https://github.com/LinYark/Sdirt.
@@ 1 Introduction
The DP sensor [15, 29] is designed to split each pixel into left and right sub-pixels, utilizing microlenses and sub-pixels to achieve phase splitting. This allows the capture of a pair of images in one shot, known as DP images, as shown in Fig.˜1(a). DP images are not only used for autofocus [31, 11, 7], but also improve performance in tasks such as deblurring [41, 18, 1] and depth estimation [13, 9, 23]. In the field of depth-from-dual-pixel (DfDP) estimation, the scarcity of DP-depth paired datasets stems from hardware limitations. To address this, many studies [37, 25, 17, 23, 3, 16] have focused on simulating DP images using RGBD datasets, which serve as more flexible and accessible data sources.
@@
Figure 1: (a) Imaging process of a DP camera. The slight shifts between the left and right DP images caused by phase differences are illustrated with white dashed lines. (b) Example comparison between real and CoC-simulated DP PSFs, showing a significant difference between them.
@@
The key to simulating DP images from RGBD datasets lies in accurately modeling the DP point spread function (PSF). In recent years, numerous model-based and calibration-based DP PSF simulators have been proposed. Calibration-based simulators [37, 16, 17] require substantial time calibrating real cameras to match DP PSFs, but suffer from issues such as interpolation errors due to discrete calibration points and the neglect of lens and DP sensor parameters, limiting model transferability. Model-based simulators [25, 23, 3] employ ideal optical or aberration analysis models to directly compute DP PSFs, bypassing calibration. However, these models often perform poorly on real DP images because they ignore lens aberrations and sensor phase splitting characteristics. As shown in Fig.˜1(b), the ideal thin-lens model’s simulated circle-of-confusion (CoC) DP PSF exhibits a significant domain gap with the real PSF, adversely affecting depth estimation.
@@
To bridge the domain gap between real DP images and DP images simulated by model-based simulators, we propose Sdirt. Sdirt consists of a ray-traced DP PSF simulator and a pixel-wise DP image rendering module, which can accurately generate DP images with aberration and phase information. Specifically, given a fixed focus lens and DP sensor system with known parameters, we use ray tracing [35, 40, 39] to calculate the spatially varying DP PSFs and train a multilayer perceptron (MLP) to predict them. Subsequently, we convolve per-pixel DP PSFs with the all-in-focus (AiF) image to simulate the DP images captured by a real camera. After training a DfDP model with DP images simulated by Sdirt, the model can accurately estimate depth using optical aberration and phase information, enhancing its ability to generalize to real DP images.
@@
We conduct extensive experiments to validate the effectiveness of Sdirt. The experimental results demonstrate that, compared to other model-based simulators, the DP PSFs and DP images generated by Sdirt better resemble real data, and the DfDP model trained on Sdirt achieves superior generalization. In summary, our contributions are fourfold:
@@
  • We propose a ray-traced DP PSF simulator that computes spatially varying DP PSFs, addressing the domain gap between simulated and real DP PSFs caused by lens aberrations and sensor phase splitting.
@@
  • Depth estimation results show that the DfDP model trained on Sdirt generalizes better to real DP images.
  • We collected DP119, a real DP-depth paired test set with an open lens structure and fixed focus, featuring diverse real-world scenes.
@@ 2.1 DP Datasets
Most professional and mobile cameras have DP sensors, but only Canon cameras [2, 26, 25] and Google Pixel phones [43, 9, 34] offer DP data. In the field of depth estimation, supervised tasks rely on DP-depth paired datasets [9, 43, 25, 2]. However, changes in the lens, sensor, focal length, or aperture settings alter the PSFs, making it necessary to re-acquire DP-depth data. Consequently, many researchers focus on simulating DP-depth datasets using RGBD data to reduce data collection costs and improve flexibility.
2.2 DP PSF simulators
Simulating DP images from RGBD datasets relies on DP PSFs, with recent progress in calibration-based [37, 17, 16] and model-based [25, 23, 3] DP PSF simulators. In calibration-based simulators, Xin et al. [37] acquired full-space DP PSFs by sampling discrete points in space and interpolating, while Li et al. [17] acquired them through size scaling using the CoC model. Furthermore, Li et al. [16] generated DP PSFs using a U-Net, which requires a large amount of DP-depth paired data. Calibration-based simulators are time-consuming and prone to interpolation and scaling errors, while model-based simulators circumvent these issues. Pan et al. [23] simplified the optical system to an ideal thin lens and used a symmetrically divided rectangular aperture to calculate DP PSFs. Punnappurath et al. [25] modeled the DP PSFs using a depletion kernel that is phase symmetric, meaning the left and right PSFs of the same object point are flip symmetric. Abuolaim et al. [3] used aberration analysis models to calculate DP PSFs, but the resulting patterns still remained phase symmetric. These studies oversimplify optical propagation by neglecting aberrations and phase splitting, leading to unrealistic domain gaps.
@@ 2.3 Applications driven by DP data
In recent years, DP data-driven applications such as depth estimation [13, 23, 9, 43, 14, 24, 10], deblurring [41, 18, 1, 38, 4, 42, 27], and refocusing [11, 7, 32, 5] have undergone rapid development. Additionally, Kang et al. [13] utilized DP data for facial scanning, similar to depth estimation tasks. Wadhwa et al. [34] employed DP data for synthetic shallow depth-of-field imaging, a popular recent application in mobile photography. Shi et al. [28] integrated DP hardware with structured diffractive optical elements for complex image encoding and high-precision multimodal reconstruction. In addition, DP data can be used for disparity estimation [36, 22], rain removal [17], and reflection removal [26]. It is foreseeable that DP data-driven applications will continue to emerge in various new fields.
Figure 2: Simulating Dual-Pixel Images from Ray Tracing pipeline. (a) Ray-traced DP PSF simulator. Calculates spatially varying DP PSFs for lens and DP sensor through ray tracing. (b) DP PSF prediction network. Trains an MLP network to predict DP PSFs, using the ray-traced DP PSFs as GT. (c) Pixel-wise DP image rendering module. The network predicts the DP PSFs for all points in the depth map (red pass). Then, each DP PSF is convolved with the AiF RGB image to render the simulated DP image (blue pass).
@@ 3 Method of Sdirt
To provide the required input for training the DfDP network, Sdirt generates realistic DP images through ray tracing. As shown in Fig.˜2(c), the module achieves pixel-wise rendering of DP images by convolving the DP PSF corresponding to each point in the depth map with the AiF RGB image. The DP PSF for each point is inferred by a pre-trained MLP network. As illustrated in Fig.˜2(b), the offline training process of the MLP network uses ray-traced DP PSFs (Fig.˜2(a)) as ground truth (GT).
Figure 3: (a) DP pixel structure layout. The left part is a right-side view, and the right part is a front perspective view. (b) On hitting the microlens, the $k$-th ray lands on left/right sub-pixel in red/blue interval, else missed. (c) Without hitting the microlens, the $k$-th ray lands on left/right sub-pixel in red/blue interval, else missed.
@@ 3.1 Ray-traced DP PSF simulator
As shown in Fig.˜2(a), lens ray tracing can accurately obtain the landing point on the sensor surface [35, 40, 39, 8]. We construct a ray set $A$ containing $n$ rays, each originating from an object point $p$. The aperture stop’s image in object space is treated as the entrance pupil, from which $n$ points are densely sampled. Each ray in $A$ is initially directed from $p$ to its corresponding entrance pupil sample. As rays in $A$ traverse each lens surface, their positions and directions are updated by refraction according to Snell’s law and the lens parameters. After multiple refractions, the rays reach the sensor plane. We denote the sensor-plane landing points as $O$ and the corresponding ray directions as $D$.
DP sensor ray tracing is limited because camera manufacturers do not disclose the structural parameters of the microlens and sub-pixel components within the DP pixel. As a result, we simplify the DP pixel structure (Fig.˜3(a)) based on past research [15, 29]. We model the microlens as a thin lens with radius $r$ and focal length $f$, and define $h$ as distance between the sub-pixel and the microlens, $w$ as sub-pixel width, and $ps$ as DP pixel size.
The DP PSF can be obtained by separately calculating the cumulative integration of the rays for each sub-pixel. We analyze which sub-pixel the $k$-th ray $A_{k}$ in the ray set $A$ ultimately enters. Through lens ray tracing, we obtain the landing point $O_{k}(x_{k},y_{k},z_{k})$ and direction $D_{k}(x_{k}^{\prime},y_{k}^{\prime},z_{k}^{\prime})$ of the $k$-th ray $A_{k}$ on the sensor surface. Based on the landing point $O_{k}$, we can easily calculate that $A_{k}$ lies within the DP pixel $(i,j)$, which has surface center coordinates $(x_{i},y_{i})$. For $O_{k}$, the situation differs depending on whether it lies within or outside the microlens:
When $O_{k}$ lies within the microlens, as shown in Fig.˜3(b), $A_{k}$ is refracted by the microlens and then enters a sub-pixel. According to geometric optics theory, $A_{k}$ and the principal ray incident on the thin lens at the same angle both converge on the focal plane of the thin lens. We only need to focus on a few boundary lines $x_{L1},x_{M1},x_{R1}$ to determine which sub-pixel $A_{k}$ ultimately enters. If $x_{k}$ is in the interval $x_{L1},x_{M1}]$, it enters the left sub-pixel; if in the interval $x_{M1},x_{R1}]$, it enters the right sub-pixel; otherwise, it is considered a missing ray. The calculation methods for the boundary lines are as follows:
@@
Where $\tan\theta$ is the tangent of the angle at which $A_{k}$ lands on the sensor surface, equivalent to ${x_{k}^{\prime}}/{z_{k}^{\prime}}$. We provide the derivation hints for Eq.˜1 in Fig.˜4(a).
When $O_{k}$ lies outside the microlens, as shown in Fig.˜3(c), $A_{k}$ enters the sub-pixel directly without refraction. Similarly, checking boundary lines $x_{L2},x_{M2},x_{R2}$ suffices to determine which sub-pixel $A_{k}$ ultimately enters:
@@
The left PSF ($PSF_{L}$) can be calculated by integrating the set of rays $A$ emitted from point $p$ over all left sub-pixels:
@@
where $\delta_{L,k}(i,j)$ represents the energy distribution of ray $A_{k}$ in the left sub-pixel at DP pixel $(i,j)$. The ray energy distribution is assumed to be a unit impulse, which equals zero unless it eventually enters the left sub-pixel of DP pixel $(i,j)$. By following the same calculation steps as for $PSF_{L}$, $PSF_{R}$ can also be computed.
Figure 4: (a) Derivation hints for Eq.˜1. (b) The valid imaging region is a frustum, and we normalize the $(x,y)$ coordinates to [1,1][-1,1]. (c) The DfDP model takes DP image as input to predict the depth map. (d) During cost volume generation in [6], we stack original disparity (green arrows) and add reverse disparity (blue arrows), with $d_{max}$ as max displacement.
@@ 3.2 Pixel-wise DP image rendering module
As shown in Fig.˜2(b), we reduce the computational cost of ray-traced DP PSFs by training an MLP network to predict them, inspired by [33, 39]. The network input is the normalized coordinate $p$ within the camera’s valid imaging region (Fig.˜4(b)), which is a frustum defined by the field of view, sensor size, and the preset minimum and maximum depths. We normalize the $(x,y)$ coordinates to [1,1][-1,1]. After fixing the focus distance, we set the DP PSF kernel size ($ks$) large enough to ensure that the DP PSF for any point within the valid imaging region can be fully displayed. The network consists of 5 hidden layers, each containing 512 neurons, and an output layer with $2*ks^{2}$ neurons. We adopt the L2 loss to supervise predictions $\hat{PSF_{L}}$ and $\hat{PSF_{R}}$:
@@
During training, we use ray-traced DP PSFs as GT and apply max normalization to alleviate learning difficulties caused by large radii PSFs. During inference, we apply sum normalization to the predicted DP PSFs, approximating the uniform intensity distribution in cameras with vignetting compensation.
Because the DP PSF is shift-variant, we convolve each pixel’s DP PSF with the AiF image to simulate the DP image captured by a camera. The RGBD dataset [30, 21] provides the paired RGB image ($I_{RGB}$) and depth map ($I_{D}$). As shown in Fig.˜2(c), we treat each pixel in $I_{D}$ as an object point and use the trained MLP to predict its DP PSF. Using $I_{RGB}$ as the AiF image, we apply pixel-wise local convolution [39] to render the DP image containing aberration and phase information.
Figure 5: Qualitative results of simulated DP PSFs. Evaluate the real and simulated (ours, CoC, L2R [3], Modeling [25], and DDDNet [23]) F/1.8 DP PSFs at two depths (0.5 m and 1.5 m) and three different positions. As the object point $p$ moves further from the optical axis, the real $PSF_{L}$ and $PSF_{R}$ become more phase asymmetric, and aberrations increase. Existing simulators neglect aberrations and DP phase splitting, causing a large gap between simulated and real DP PSFs. Only our ray-traced simulator predicts realistic results at all depths and positions.